OpenStudy (babynini):

Write the expression as a product and simplify sin2alpha - sin8 alpha. Will give medal!

OpenStudy (babynini):

@optiquest

OpenStudy (anonymous):

Sum to Product Formulas http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (babynini):

Which one is it?

OpenStudy (babynini):

2nd one yeah?

OpenStudy (anonymous):

Sum to Product Formula

OpenStudy (anonymous):

yea sin - sin

OpenStudy (babynini):

em..in this case alpha = 2 beta = 8?

OpenStudy (babynini):

or does alpha = 2alpha beta = 8 alpha ??

OpenStudy (anonymous):

everything inside the first sin in alpha

OpenStudy (anonymous):

is

OpenStudy (babynini):

2cos[(alpha+beta)/2]sin[(alpha-beta)/2] would be 2cos[(2alpha + 8alpha)/2]sin[(2alpha - 8alpha)/2] ??

OpenStudy (anonymous):

yea

OpenStudy (babynini):

um so in the parenthesis when its ((2alpha + 8 beta)/2) that would become 5alpha?

OpenStudy (babynini):

and sin((2alpha-8alpha)/2 = -3alpha ?

OpenStudy (anonymous):

yes

OpenStudy (babynini):

So we have 2cos(5alpha)sin(-3alpha) now what?

OpenStudy (anonymous):

should be it

OpenStudy (babynini):

what? really? o.0

OpenStudy (babynini):

it can't be simplified further?

OpenStudy (anonymous):

dont think so

OpenStudy (babynini):

what about double angle formulas? those don't work here?

OpenStudy (anonymous):

try it and see

OpenStudy (babynini):

I don't know, that's why i'm asking :P double angle formula = sin(2theta) = 2sin (theta) cos(theta) what we have is 2cos(5alpha)sin(-3alpha) fit or no?

OpenStudy (anonymous):

think it needs 2 theta like the name

OpenStudy (anonymous):

i dont remember the derivation but i dont think it applies here

OpenStudy (babynini):

okies. Thank you!

OpenStudy (babynini):

@zepdrix dude..I know you're probably tired of me now. but can you just check reaaal quick if this is the most it can be simplified?

OpenStudy (babynini):

I will love you forever! and (maybe..not) stop bugging you!

zepdrix (zepdrix):

I guess you could apply one more tiny tiny step if you wanted :) Sine is an `odd function` therefore \(\Large\rm \sin(-x)=-\sin(x)\) Do you see how we can use that in our problem here?\[\Large\rm 2\sin(-3\alpha)\cos(5\alpha)\]

OpenStudy (babynini):

-2sin(3alpha)cos(5alpha)

zepdrix (zepdrix):

Good \c:/ That's probably as far as you need to simplify it. If we had even multiples of alpha, then yes we could go further and apply our sine and cosine double angle formulas. But I don't think they want us to do that here. That's a good place to stop!

OpenStudy (babynini):

yay! thanks :)